# Unifying KV Cache Compression for Large Language Models with LeanKV

### v2

@@ Abstract
Large language models (LLMs) exhibit exceptional performance but incur significant serving costs due to their substantial memory requirements, with the key-value (KV) cache being a primary bottleneck. Existing KV cache compression techniques, such as quantization and pruning, apply uniform treatment to both keys and values, and discard unimportant tokens entirely, overlooking the fine-grained differences in significance of various components within the KV cache. To address these limitations, we introduce LeanKV, a framework that advances KV cache compression by exploiting three levels of differentiation in the KV cache: (1) the differing impact of keys and values on attention computation, (2) the varying importance of tokens, and (3) the diverse dynamic sparsity patterns across attention heads. At the core of LeanKV is an on-GPU memory manager that compacts fragmented free memory list into contiguous regions in parallel, effectively translating sparsity in the KV cache into performance gains. We evaluate LeanKV on several mainstream models, including the recent "thinking model". LeanKV is able to compress the KV cache by 2.7×2.7$\times$2.7 × to 5.7×5.7$\times$5.7 × with near-lossless accuracy on complex workloads requiring sophisticated reasoning and long-generation capabilities, and enhances throughput by 1.9×1.9$\times$1.9 × to 5.4×5.4$\times$5.4 ×.
@@ 1 Introduction
LLMs are fundamentally based on the Transformer architecture [61, 62], which employs the attention mechanism to generate tokens sequentially, based on the input prompt and previously generated tokens. Within the attention mechanism, each token is transformed into query, key and value feature vectors. The attention scores, derived from the dot product of the query and key vectors, enable the model to weigh the significance of different tokens. Subsequently, the output is computed as the weighted sum of the value vectors, based on the attention scores. To capture diverse relationships between tokens, the attention mechanism is typically divided into multiple heads [61, 51, 5]. Each head independently computes attention, and their outputs are concatenated and transformed to create a richer representation of the input data.
@@
To avoid redundant computation, LLM inference frameworks typically cache intermediate key and value tensors in memory, commonly referred to as the KV cache [32]. The size of the KV cache scales linearly with both sequence length and the number of concurrent requests, often comprising over 90% of total memory consumption [32, 67, 19, 66]. As state-of-the-art models continue to support longer sequences [17, 30, 38, 4, 57, 60], and with the rise of recent "thinking models" [23, 29, 58] that generate extended reasoning processes, the KV cache has emerged as a critical bottleneck for LLM serving efficiency. It limits the number of concurrent requests that can fit in memory and increases attention computation latency due to its memory bandwidth-bound nature [15, 14].
Recognizing the critical role of the KV cache in limiting serving efficiency, researchers have investigated various compression techniques, primarily focused on quantization [37, 2, 27] and pruning [69, 64, 10, 20, 35]. Quantization reduces the size of the KV cache by converting floating-point feature values into integers, representing them based on pre-defined intervals. Pruning, on the other hand, reduces the token sequence length by eliminating unimportant tokens based on attention scores. While promising, these techniques overlook the fine-grained differentiation in the significance of various KV cache components, limiting their overall effectiveness. First and foremost, existing quantization methods save key and value vectors of all tokens at a uniform precision, disregarding the distinct roles of keys and values and the varying importance of tokens. Second, current pruning methods employ static memory allocation across attention heads, neglecting the dynamic and heterogeneous attention patterns observed across heads and requests. Moreover, these pruning methods have not been comprehensively evaluated on workloads requiring advanced reasoning, such as programming and mathematics.
In this paper, we present LeanKV, a framework that advances KV cache compression by leveraging three levels of differentiation within the KV cache:
@@
1. Differentiated Keys and Values. LeanKV assigns higher precision to key vectors than to value vectors, based on the observation that key vectors play a more significant role in attention computations.
@@
2. Token Importance Differentiation. Guided by attention scores, LeanKV stores tokens at varying precision levels based on their significance. Most critical tokens are quantized at higher precision, less critical tokens are quantized at lower precision, and the least significant tokens are pruned. This hierarchical approach enables a flexible tradeoff between model accuracy and memory efficiency.
@@
3. Per-head Dynamic Sparsity. Attention score distributions exhibit dynamic sparsity across both heads and requests. Within a single request, the number of critical tokens varies across heads, and for the same head, the number of critical tokens also varies across different requests. To address this variability, LeanKV dynamically identifies critical tokens per-head, per-request, adapting memory allocation accordingly.
These differentiations wthin the KV cache pose new challenges in memory management. In LeanKV, each request requires varying amounts of memory for each attention head to store tokens at different precision levels. Moreover, key and value vectors of the same token are compressed at different precisions. These complexities significantly increase memory management overhead, compared to traditional attention mechanisms, where memory allocation is uniform across all heads.
@@
To address these challenges, LeanKV employs an on-GPU memory manager that efficiently handles the irregular memory usage patterns across heads and requests via parallel KV compaction. Rather than performing memory management operations sequentially on CPU for each head and request, parallel KV compaction optimizes memory allocation and recycling by packing fragmented free memory lists into contiguous regions directly on the GPU in parallel. Parallel KV compaction is supported by three key on-GPU data structures:
@@
1. Unified Pages. GPU memory is divided into fixed-size pages containing vectors stored at various precisions. Each page is dynamically parsed according to the precision requirements of the head and request, while an attention GPU kernel is co-designed with the page layout to maximize memory bandwidth utilization.
@@
2. Circular Free Page List. All page IDs are managed in a centralized, GPU-resident circular list. Both free and used pages are kept in contiguous regions, enabling efficient page allocation and recycling via parallel prefix sum.
@@
3. Bidirectional Page Table. Instead of maintaining separate page tables for high- and low-precision pages, LeanKV consolidates both into a single bi-directional page table to minimize metadata overhead. High precision page IDs grow from the left side of the table, and low-precision page IDs grow from the right side.
We implement LeanKV on vLLM [32], a state-of-the-art LLM inference system, and evaluate it across three prominent model families including the emerging thinking model. The evaluation is conducted on representative workloads that demand advanced reasoning capabilities, such as mathematics and coding. Our experimental results demonstrate that LeanKV offers a superior cost-accuracy tradeoff compared to existing KV cache management methods. Specifically, LeanKV compresses the KV cache by 2.7×2.7$\times$2.7 × to 5.7×5.7$\times$5.7 × with near-lossless accuracy, resulting in throughput improvement of 1.9×1.9$\times$1.9 × to 5.4×5.4$\times$5.4 ×. The throughput improvement is primarily enabled by the parallel KV compaction mechanism, which reduces management overhead by three orders of magnitude, decreasing the proportion of memory management time in total execution latency from up to $76\%$ to less than $1\%$ compared to a sequential CPU-based implementation.
@@ 2 Background and Related Work
In this section, we describe Transformer-based large language models and review existing techniques for KV cache management aimed at improving LLM serving efficiency.
@@ 2.1 Large Language Models
Transformer Architecture. Large language models (LLMs) are predominantly built upon the Transformer architecture [61], which is used to capture complex long-range dependencies in sequential data. At the core of Transformer architecture is the attention mechanism, which allows each token in a sequence to weigh the significance of other tokens when constructing its contextualized representation. Mathematically, the attention computation is defined as:
@@
$\mathbf{Q}$ and $\mathbf{K}$ are matrices of size $l\times d$, representing the queries and keys, respectively, where $l$ denotes the number of tokens and $d$ represents the feature dimensionality. Vectors $\mathbf{q}_{i}$, $\mathbf{k}_{i}$, and $\mathbf{v}_{i}$ correspond to the query, key, and value of the $i^{\text{th}}$ token. The dot product between $\mathbf{Q}$ and $\mathbf{K}$ measures the relevance between each pair of query and key. The softmax function normalizes these dot products to form a probability distribution, referred to as attention scores, which are used to compute a weighted sum of the value vectors and produce the output.
To capture a broader range of interactions between tokens, Transformer models employ multi-head attention (MHA), where each head independently computes attention using distinct projections of queries, keys, and values. The outputs from different heads are concatenated and linearly transformed to create a more expressive representation. Grouped-query attention (GQA) [5] improves the efficiency of MHA by allowing a group of query heads to share the same projection of keys and values, referred to as a KV head.
Autoregressive Generation and KV Cache. The autoregressive generation process of LLMs consists of two phases: the prompt phase, during which the model computes latent representations for all tokens in the prompt, the user-provided context, and generates the first new token, and the generation phase, where the model iteratively generates subsequent tokens by attending to both the prompt and previously generated tokens. To avoid redundant computations across generation steps, KV cache is introduced to store the keys and values of all previous tokens. However, the size of the KV cache grows linearly with both the sequence length and batch size, quickly becoming a bottleneck for inference throughput [67, 32]. Therefore, efficient KV cache management is critical for alleviating memory bottlenecks and improving LLM serving efficiency.
@@ 2.2 KV Cache Optimization
KV Cache Quantization. Quantization [43, 16, 22, 27] reduce the KV cache size by approximating high-precision floating points with discrete low-bit values. To quantize a given tensor $\mathbf{X}$, we first calculate the scale $s$ and zero point $z$ as shown in Eq. (2), where $\mathbf{X}_{\text{max}}$ and $\mathbf{X}_{\text{min}}$ represent the maximum and minimum values in $\mathbf{X}$, and $b$ is the bit-width. Then, quantization is applied element-wise as: $\mathbf{Q}=\operatorname{round}\left(\frac{\mathbf{X}+z}{s}\right)$. During inference, the original values are approximately reconstructed via dequantization: $\hat{\mathbf{X}}=s\cdot\mathbf{Q}-z$. Unlike the quantized tensor $\mathbf{Q}$, the metadata $s$ and $z$ are kept in higher precision (e.g., FP16) to ensure more accurate dequantization.
$s=\frac{\mathbf{X}_{\text{max}}-\mathbf{X}_{\text{min}}}{2^{b}-1},\quad z=-\mathbf{X}_{\text{min}}$ (2)
State-of-the-art KV cache quantization methods such as Atom [70] and Qserve [37] apply this process to each key and value vector independently. However, these uniform quantization methods, which apply a fixed bit-width across keys and values of all tokens, may not be optimal. They overlook the varying importance of tokens as well as the different roles of keys and values in the attention calculation.
KV Cache Pruning. Several recent works [69, 64, 35, 10, 40, 20] have explored KV cache pruning to alleviate the memory bottleneck in LLM inference, which can be considered as an extreme case of quantization. These methods use attention scores to assess token importance and evict less important tokens, thereby reducing the KV cache size. The effectiveness of these methods, however, is limited by their static and inflexible allocation of memory resources. Specifically, H2O [69] and SnapKV [35] allocate the same memory budget uniformly across all attention heads and layers. PyramidKV [10] allocates more memory in lower layers and less in higher ones, based on the empirical observation that the number of important tokens shrinks as the model depth increases, but still relies on static heuristics and cannot dynamically adjust cache allocations to suit various workloads. Complementary to KV cache pruning, another line of work [52, 33] addresses GPU memory constraints by offloading the entire KV cache to CPU memory and selectively loading important tokens to the GPU for attention computations. However, these methods do not inherently reduce the KV cache size and incur significant latency due to data transfers between the CPU and GPU.
@@ 3 The Case for Differentiated KV
We present the case for three levels of differentiation within the KV cache, motivating the design of LeanKV.
@@ 3.1 Differentiated Impacts of Keys and Values
Beyond the broader impact of key vectors, it is crucial to establish the relative significance of attention scores, determined by key vectors, over value vectors. Without this, the significance of key vectors could be overshadowed by that of value vectors. To this end, we reformulate the attention mechanism as a weighted sum of unit vectors, as presented in Equation 3. This formulation decomposes each input token’s contribution into two components: the unit vector $\frac{v_{j}}{|v_{j}|}$, representing only the direction of the token in feature space, and a coefficient, defined as the product of the attention score and the norm of the value vector, which determines the relative significance of the token’s direction. This allows us to assess the relative importance of attention scores and value vectors by examining their contributions to these coefficients.
@@
Figure 1 presents the distributions of the average attention score per token and the norms of value vectors across three representative layers of the Llama3-8B model [17], evaluated on the first one thousand articles from the Wikitext dataset [42]. Notably, the attention scores span seven orders of magnitude, vastly exceeding the range of value vector norms, which only cover at most two orders of magnitude. Similar patterns are also observed in the larger Llama3-70B model. This pronounced disparity highlights the pivotal role of attention scores in determining each token’s contribution to the attention output. We therefore conclude that key vectors exert a broader and more impactful influence than value vectors, motivating further exploration of processing key and value vectors at distinct precision levels.

### v3

# DiffKV: Differentiated Memory Management for Large Language Models with Parallel KV Compaction

@@ Abstract
Large language models (LLMs) demonstrate remarkable capabilities but face substantial serving costs due to their high memory demands, with the key-value (KV) cache being a primary bottleneck. State-of-the-art KV cache compression techniques, such as quantization and pruning, apply uniform treatment to both keys and values, and discard unimportant tokens entirely, overlooking the fine-grained distinctions in the significance of individual KV cache components. To address such limitations, we introduce DiffKV, a novel framework for efficient KV cache compression that exploits three levels of differentiation in the KV cache: (1) the differing impact of keys and values on attention computation, (2) the varying importance of tokens, and (3) the diverse dynamic sparsity patterns across attention heads. These levels of differentiation introduce irregular memory usage patterns across different requests and attention heads, posing significant scalability challenges for memory management. To address these challenges, DiffKV proposes an on-GPU memory manager that compacts fragmented free memory list into contiguous regions in parallel, effectively translating sparsity in the KV cache into performance gains. We evaluate DiffKV on several mainstream LLMs, including the emerging thinking models that generate extended chains of thought. DiffKV is able to compress the KV cache by $2.7\times$ to $5.7\times$ with near-lossless accuracy on complex workloads requiring sophisticated reasoning and long-generation capabilities, and enhances throughput by $1.9\times$ to $5.4\times$. Source codes of DiffKV are available at https://github.com/zyqCSL/DiffKV.
@@ 1 Introduction
To avoid redundant computation, LLM inference frameworks typically cache intermediate key and value tensors in memory, commonly referred to as the KV cache [39]. The size of the KV cache scales linearly with both sequence length and the number of concurrent requests, often comprising over 90% of total memory consumption [39, 74, 22, 73]. As state-of-the-art models continue to support longer sequences [20, 36, 45, 4, 66, 68], and with the rise of recent thinking models [26, 35, 67] that generate extended reasoning processes, the KV cache has emerged as a critical bottleneck for LLM serving efficiency. It limits the number of concurrent requests and increases attention computation latency due to its memory bandwidth bound nature [17, 16].
Researchers have investigated various compression techniques, primarily focused on pruning [76, 71, 11, 23, 42] and quantization [44, 2, 32]. Pruning reduces the token sequence length by eliminating unimportant tokens based on attention scores. Quantization, on the other hand, reduces the size of the KV cache by converting floating-point feature values into lower-precision representations.
While effective, these methods overlook the nuanced variations in importance across different components of the KV cache. First, existing pruning approaches employ a one-size-fits-all strategy, statically allocating memory uniformly across attention heads despite the per-head dynamic attention sparsity patterns. Second, current quantization schemes apply uniform precision to all key and value vectors, ignoring both the distinct roles of keys versus values in the attention mechanism, and the varying significance of different tokens. Finally, both pruning and quantization treat all requests uniformly, overlooking variability in information density across requests.
To address these limitations and advance KV cache compression, we propose DiffKV, a novel framework that exploits three levels of differentiation within the KV cache:
@@
1. Differentiated Keys and Values. DiffKV assigns higher precision to key vectors than to value vectors, based on the observation that keys (critical for attention score computation) play a more significant role than values (used primarily for weighted aggregation) in attention computations.
@@
2. Token Importance Differentiation. Guided by attention scores, DiffKV stores tokens at varying precision levels based on their significance. Most critical tokens are quantized at higher precision, less critical tokens are quantized at lower precision, and the least significant tokens are pruned.
@@
3. Per-head Dynamic Sparsity. Attention score distributions exhibit dynamic sparsity across both heads and requests: the number of important tokens varies across different heads, and for the same head under different requests. To address this variability, DiffKV dynamically identifies critical tokens per-head, per-request, adapting memory allocation accordingly.
Figure 1: KV cache memory allocation patterns for (a) pruning, (b) uniform quantization, and (c) DiffKV across two attention heads in two 5-token requests. Boxes represent retained tokens (annotated with token IDs), with box size proportional to memory usage. Percentages indicate total memory usage relative to an uncompressed KV cache.
Figure 1 compares the KV cache usage of pruning, quantization, and DiffKV across two heads in two 5-token requests. In both requests, pruning uniformly eliminates one token per head (reducing cache usage to 80%), failing to adapt to heads with naturally sparser attention patterns. Quantization, meanwhile, applies INT8 precision uniformly to all FP16 values, halving memory usage but missing opportunities for finger-grained savings by treating keys and values differently based on their relative importance. Moreover, both approaches treat two requests identically, overlooking their distinct sparsity patterns.
In contrast, DiffKV exhibits more adaptive memory usage. In Request 1, for Head A, DiffKV prunes one token and further distinguishes precision across the rest: Token 0, deemed highly important, is stored at high precision (key in INT8, value in INT4), while Tokens 1, 3, and 4 are stored at lower precision (keys in INT4, values in INT2). This reduces usage to 25% for keys and 12.5% for values. For Head B, DiffKV instead identifies two tokens as highly important, resulting in 30% and 15% KV cache usage for keys and values, respectively. Averaged across both heads, DiffKV uses 20.6% of memory for Request 1, outperforming pruning and quantization. Additionally, DiffKV also adapts flexibly across requests. In Request 2, Head A stores one token at high precision and the remaining four at lower precision without pruning, whereas Head B prunes two tokens and stores two others at high precision. These examples highlight DiffKV’s ability to leverage both per-head dynamic sparsity and request-level variation, enabling more aggressive and flexible compression than prior methods.
Although DiffKV achieves a high KV cache compression ratio, it introduces significant memory usage irregularity. As depicted in Figure 1, pruning and quantization maintain uniform KV cache utilization across attention heads. In contrast, DiffKV exhibits substantial variability, not only across different heads but also between the key and value within individual heads. To mitigate fragmentation, the memory manager must precisely track and allocate memory per head and per request. However, this introduces a scalability challenge: even a modestly sized LLM like Llama3-8B [20] comprises hundreds of attention heads, and with hundreds of concurrent requests, the memory manager must handle tens of thousands of heterogeneous memory regions during each inference step. Considering that model execution time per step is on the order of tens of milliseconds, improper memory management could render the overhead of dynamic allocation prohibitive, negating the benefits of KV cache compression.
@@
To overcome the scalability challenge, DiffKV introduces an on-GPU memory manager that efficiently handles the irregular memory usage patterns via parallel KV compaction, which optimizes memory allocation and recycling by packing fragmented free memory lists into contiguous regions directly on the GPU in parallel. Parallel KV compaction is enabled by three essential on-GPU data structures:
@@
1. Unified Pages. GPU memory is partitioned into fixed-size pages, each storing tokens at a specific precision. For example, a page may hold tokens with keys in INT8 and values in INT4. Each page is dynamically configured and parsed according to the precision requirements of the head and request.
@@
2. Circular Free Page List. All page IDs are managed in a centralized, GPU-resident circular list. Both free and used pages are kept in contiguous regions, enabling efficient page allocation and recycling via parallel prefix sum [29].
@@
3. Bidirectional Page Table. Instead of maintaining separate page tables for high- and low-precision pages, DiffKV consolidates both into a single bi-directional page table to minimize metadata overhead. High-precision page IDs grow from the left side of the table, and low-precision page IDs grow from the right side. Additional precision levels can also be efficiently accommodated by employing multiple bi-directional page tables.
We have implemented DiffKV on vLLM [39] and evaluate it across multiple LLMs, including the emerging thinking models QwQ-32B [67], R1-Distill-Qwen-14B and R1-Distill-Llama-8B [26]. Our experiment results demonstrate that Diff-
KV offers a superior cost-accuracy tradeoff. Specifically, Diff-
KV compresses the KV cache by $2.7\times$ to $5.7\times$ with near-lossless accuracy, resulting in throughput improvement of $1.9\times$ to $5.4\times$. Notably, DiffKV is, to the best of our knowledge, the first KV cache compression framework evaluated on thinking models and complex reasoning tasks that require advanced chain-of-thought (CoT) capabilities, while achieving FP16-comparable generation quality.
@@ 2 Background and Related Work
In this section, we describe Transformer-based large language models and review existing techniques for KV cache management.
@@ 2.1 Large Language Models
Large language models (LLMs) are predominantly built upon the Transformer architecture [69]. At the core of Transformer is the attention mechanism, which allows each token in a sequence to weigh the significance of other tokens when constructing its contextualized representation. During autoregressive inference, the attention mechanism operates in a causal manner, attending only to preceding tokens. Mathematically, the standard attention computation is defined as:
@@
$\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ are matrices of size $l\times d$, representing the queries, keys and values respectively, where $l$ denotes the number of tokens processed so far in the sequence and $d$ represents the feature dimensionality. Vectors $\mathbf{q}_{i}$, $\mathbf{k}_{i}$, and $\mathbf{v}_{i}$ correspond to the query, key, and value of the $i^{\text{th}}$ token.
To capture a broader range of interactions between tokens, Transformer models employ multi-head attention (MHA), where each head independently computes attention using distinct projections of queries, keys, and values. Grouped-query attention (GQA) [6, 60] improves the efficiency of MHA by allowing a group of query heads to share the same projection of keys and values, referred to as a KV head.
LLM execution involves two phases: the prompt phase, where the model computes latent representations for all tokens in the prompt and generates the first new token, and the generation phase, where subsequent tokens are generated iteratively. To avoid redundant computations across generation steps, KV cache is introduced to store the keys and values of all previous tokens. However, the size of the KV cache grows linearly with both the sequence length and batch size, quickly becoming a bottleneck for inference throughput [74, 39]. Therefore, efficient KV cache management is critical for alleviating memory bottlenecks and improving LLM serving efficiency.
@@ 2.2 KV Cache Optimization
KV Cache Quantization. Quantization [50, 18, 25, 32] reduces the KV cache size by approximating high-precision floating points with discrete low-bit values. For a tensor $\mathbf{X}$, we first compute the scale $s$ and zero point $z$ based on $\mathbf{X}_{min}$ and $\mathbf{X}_{max}$, then apply quantization element-wise as: $\mathbf{Q}=\operatorname{round}\left(\frac{\mathbf{X}-z}{s}\right)$. During inference, the original tensor is approximately reconstructed via dequantization: $\hat{\mathbf{X}}=s\cdot\mathbf{Q}+z$. Unlike the quantized tensor $\mathbf{Q}$, the metadata $s$ and $z$ are kept in higher precision (e.g., FP16) to ensure more accurate dequantization. State-of-the-art KV cache quantization methods such as Atom [77] and Qserve [44] apply this process to each key and value vector independently. However, these uniform quantization methods, which apply a fixed bit-width across keys and values of all tokens, may not be optimal. They overlook the varying importance of tokens as well as the different roles of keys and values in the attention calculation.
KV Cache Pruning. Several recent works [76, 71, 42, 11, 47, 23] have explored KV cache pruning to alleviate the memory bottleneck in LLM inference, which can be considered as an extreme case of quantization. These methods use attention scores to assess token importance and evict less important tokens, thereby reducing the KV cache size. The effectiveness of these methods, however, is limited by their static and inflexible allocation of memory resources. Specifically, H2O [76] and SnapKV [42] allocate the same memory budget uniformly across all attention heads and layers. PyramidKV [11] allocates more memory in lower layers and less in higher ones, based on the empirical observation that the number of important tokens shrinks as the model depth increases, but still relies on static heuristics and cannot dynamically adjust cache allocations to suit various workloads. Complementary to KV cache pruning, another line of work [61, 40] addresses GPU memory constraints by offloading the entire KV cache to CPU memory and selectively loading important tokens to the GPU for attention computations. However, these methods do not inherently reduce the KV cache size and incur additional latency due to data transfers between the CPU and GPU.
@@ 3 The Case for Differentiated KV
We motivate the design of DiffKV by presenting the case for three levels of differentiation within the KV cache: (i) the differentiated impacts of keys and values (Section 3.1), (ii) differentiated token importance (Section 3.2), and (iii) per-head dynamic sparsity (Section 3.3). Building on these insights, Section 3.4 describes how DiffKV combines them together and highlights its contributions over prior systems.
@@ 3.1 Differentiated Impacts of Keys and Values
Beyond the broader impact of key vectors, it is crucial to establish the relative significance of attention scores, determined by key vectors, over value vectors. Without this, the significance of key vectors could be overshadowed by that of value vectors. To this end, we reformulate the attention mechanism as a weighted sum of unit vectors, as presented in Equation 2. This formulation decomposes each input token’s contribution into two components: the unit vector $\frac{v_{j}}{|v_{j}|}$, obtained by dividing the token’s value vector by its L2 norm to capture only its direction in feature space, and a coefficient, defined as the product of the attention score and the norm of the value vector, which determines the relative significance of the token’s direction. This allows us to assess the relative importance of attention scores and value vectors by examining their contributions to these coefficients.
@@
Figure 2 presents the distributions of the average attention score per token and the norms of value vectors across three representative layers of the Llama3-8B model [20], evaluated on the first one thousand articles from the Wikitext dataset [49]. Notably, the attention scores span seven orders of magnitude, vastly exceeding the range of value vector norms, which only cover at most two orders of magnitude. Similar patterns are also observed in the larger Llama3-70B model. This pronounced disparity highlights the pivotal role of attention scores in determining each token’s contribution to the attention output. We therefore conclude that key vectors exert a broader and more impactful influence than value vectors, motivating further exploration of processing key and value vectors at distinct precision levels. We validate this intuition in Section 7.2, showing that K8V4 and K4V2, which quantize keys to 8/4 bits and values to 4/2 bits respectively, outperform their mirror configurations K4V8 and K2V4.